OpenStudy (anonymous):

@jim_thompson5910 I need a lot of help can you be the hero for me today

OpenStudy (anonymous):

lol im your hero wassup ... medl to opperate

jimthompson5910 (jim_thompson5910):

how many problems?

OpenStudy (anonymous):

umm 9

jimthompson5910 (jim_thompson5910):

we'll see how it goes, but 9 sounds a bit too high

jimthompson5910 (jim_thompson5910):

but let's do a few

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\sqrt{(+3)^2)^2}\]

OpenStudy (anonymous):

I have to simplify it

jimthompson5910 (jim_thompson5910):

seems like a typo...just +3 ?

OpenStudy (anonymous):

\[\sqrt{(x+3)^2}\] sorry

jimthompson5910 (jim_thompson5910):

ok from last night, I was trying to get you to see that square roots undo squaring

jimthompson5910 (jim_thompson5910):

basically, you take any positive number and you square it then if you take the square root of that result, you get back to the original positive number

jimthompson5910 (jim_thompson5910):

example 5^2 = 25 sqrt(25) = 5

jimthompson5910 (jim_thompson5910):

so that's why sqrt(x^2) = x when x is some positive number

OpenStudy (anonymous):

okay so I square 3

jimthompson5910 (jim_thompson5910):

so if you take some unknown number x+3 and you square it, you get (x+3)^2

jimthompson5910 (jim_thompson5910):

then take the square root of that to get ???

OpenStudy (anonymous):

that is where I get lost at

jimthompson5910 (jim_thompson5910):

do you see how sqrt(x^2) = x

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

so think of x+3 as a number (just like x is)

jimthompson5910 (jim_thompson5910):

if you want, let z = x+3 so sqrt( (x+3)^2 ) is the same as sqrt( z^2 )

OpenStudy (anonymous):

okay that is 4 right

jimthompson5910 (jim_thompson5910):

idk how you're getting 4

OpenStudy (anonymous):

ohh sorry

jimthompson5910 (jim_thompson5910):

what does sqrt( z^2 ) simplify to

OpenStudy (anonymous):

z^2

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

square roots undo squaring

OpenStudy (anonymous):

okay so wat would it be

jimthompson5910 (jim_thompson5910):

Are you sure you understand how \[\Large \sqrt{x^2} = x\] when x is positive

OpenStudy (anonymous):

I DONT THINK SO

jimthompson5910 (jim_thompson5910):

ok replace x in that equation with any number you want

jimthompson5910 (jim_thompson5910):

pick any positive number and tell me which one you pick